OpenStudy (anonymous):

Solve the equation 5^(x-1)=3^x. Round the answer to 3 decimal places.

OpenStudy (anonymous):

Take a log of both sides to move the variables out of the exponents.

OpenStudy (anonymous):

(x-1)log5=xlog3 (x-1)/x=log3/log5 (x-1)/x= 0.68261 x-1 = 0.68261x -1=-.31739x 3.151